# Self-Dual Codes and LCD Codes in Sum-Rank Metric

### v1

@@ 5 Asymptotically Good Self-Dual Sum-Rank Codes
In this section, we aim to prove that there exists an asymptotically good sequence of self-dual sum-rank codes. In order to achieve the goal, we will use the second method above. In the proof of the following theorem, we will use an important theorem called Bolzano-Weierstrass Theorem which can be found in [65]. It claims that for any bounded sequence of real numbers, there exists a convergent subsequence of it. What is more, it is necessary to recall that $\mu_{q}(n)$ is the minimal size of the non-trivial $q$-cyclotomic cosets.
@@
As $q$ is even, then $-1=1$ in $\mathbb{F}_{q^{2}}$. Thus, $-1$ is a square in $\mathbb{F}_{q^{2}}$. Since $t_{1}$, $t_{2}$, $\ldots$ are odd positive integers satisfying that $t_{i}$ goes to infinity with $i$ and $\lim\limits_{i\rightarrow\infty}\frac{\log_{q^{2}}t_{i}}{\mu_{q^{2}}(t_{i})}=0$, $t_{1}$, $t_{2}$, $\ldots$ are odd positive integers coprime to $q^{2}$ satisfying that $t_{i}$ goes to infinity with $i$ and $\lim\limits_{i\rightarrow\infty}\frac{\log_{q^{2}}t_{i}}{\mu_{q^{2}}(t_{i})}=0$. Hence, by Lemma 2.3, there exist Euclidean self-dual codes $C^{\prime}_{i}\subseteq\mathbb{F}_{q^{2}}^{2t_{i}}$, $i=1,2,\ldots$ such that the code sequence $C_{1}^{\prime}$, $C_{2}^{\prime}$, $\ldots$ is asymptotically good.
@@
As a self-dual basis $\mathcal{B}$ of $\mathbb{F}_{q}^{m}$ over $\mathbb{F}_{q}$ exists if and only if $q$ is even or both $q$ and $m$ are odd, there is no self-dual basis of $\mathbb{F}_{q^{2}}$ over $\mathbb{F}_{q}$ if $q$ is power of an odd prime. Thus, in this case, even if $C$ is self-dual, we cannot guarantee that $Mat_{\mathcal{B}}(C)$ is a self-dual sum-rank code. This is the reason why we restrict $q$ to be even.
@@ 6 Conclusion
In this paper, we introduce the notions of self-dual sum-rank codes and LCD sum-rank codes and give some properties of self-dual sum-rank codes. Then we connect them with Euclidean self-dual sum-rank codes and LCD sum-rank codes in two ways and provide two methods of constructing them by using the two connections respectively. Using these methods, We construct not only some normal self-dual codes and LCD codes in sum-rank metric with good parameters but also some cyclic self-dual sum-rank codes and cyclic LCD sum-rank codes. We also prove that asymptotically good self-dual sum-rank codes exist by the second method.
@@
    Let $C_{i}$ be a linear code in $\mathbb{F}_{q^{m}}^{t}$, $10\leq i\leq m-$ 1. Is it true that $SR(C_{0},C_{1},\ldots,C_{m-1})^{\bot_{tr}}=SR(C_{0}^{\bot},C_{1}^{\bot},\ldots,C_{m-1}^{\bot})$?
In addition, giving more constructions of self-dual sum-rank codes and LCD sum-rank codes with good parameters will be a significant direction. What is more, it is also important to compute the parameters of those known self-dual sum-rank codes and LCD sum-rank codes.
@@
$1$ $1+x$ $2$ $d_{sr}(Mat_{\mathcal{B}}(C))=1$
$2$ $1+x^{2}$ $2$ $d_{sr}(Mat_{\mathcal{B}}(C))=2$
$3$ $w^{2}+w^{2}x+x^{2}+x^{3}$, $w+wx+x^{2}+x^{3}$ $3$ $d_{sr}(Mat_{\mathcal{B}}(C))=2$
$4$ $1+x^{4}$ $2$ $d_{sr}(Mat_{\mathcal{B}}(C))=2$
$5$ $1+x^{5}$ $2$ $d_{sr}(Mat_{\mathcal{B}}(C))=2$
$6$ $w^{2}+w^{2}x+x^{2}+wx^{3}+w^{2}x^{4}+x^{5}+x^{6}$, $w+wx+x^{2}+w^{2}x^{3}+wx^{4}+x^{5}+x^{6}$ $4$ $2\leq d_{sr}(Mat_{\mathcal{B}}(C))\leq 4$
$7$ $1+x+x^{2}+x^{3}+x^{6}+x^{7}$, $1+x+x^{4}+x^{5}+x^{6}+x^{7}$ $4$ $2\leq d_{sr}(Mat_{\mathcal{B}}(C))\leq 3$
$8$ $1+x^{8}$ $2$ $d_{sr}(Mat_{\mathcal{B}}(C))=2$
$9$ $w^{2}+w^{2}x+x^{2}+wx^{3}+w^{2}x^{4}+x^{5}+wx^{6}+w^{2}x^{7}+x^{8}+x^{9}$, $w+wx+x^{2}+w^{2}x^{3}+wx^{4}+x^{5}+w^{2}x^{6}+wx^{7}+x^{8}+x^{9}$ $4$ $2\leq d_{sr}(Mat_{\mathcal{B}}(C))\leq 4$
$10$ $1+x^{10}$ $2$ $d_{sr}(Mat_{\mathcal{B}}(C))=2$
$11$ $1+x+wx^{2}+wx^{3}+x^{4}+x^{5}+x^{6}+x^{7}+w^{2}x^{8}+w^{2}x^{9}+x^{10}+x^{11}$, $1+x+w^{2}x^{2}+w^{2}x^{3}+x^{4}+x^{5}+x^{6}+x^{7}+wx^{8}+wx^{9}+x^{10}+x^{11}$ $6$ $3\leq d_{sr}(Mat_{\mathcal{B}}(C))\leq 6$
$12$ $1+x+w^{2}x^{2}+wx^{3}+wx^{5}+x^{6}+w^{2}x^{7}+w^{2}x^{9}+wx^{10}+x^{11}+x^{12}$, $w+wx^{2}+x^{4}+w^{2}x^{6}+wx^{8}+x^{10}+x^{12}$, $w^{2}+w^{2}x+x^{2}+wx^{3}+w^{2}x^{4}+x^{5}+wx^{6}+w^{2}x^{7}+x^{8}+wx^{9}+w^{2}x^{10}+x^{11}+x^{12}$, $w+wx+x^{2}+w^{2}x^{3}+wx^{4}+x^{5}+w^{2}x^{6}+wx^{7}+x^{8}+w^{2}x^{9}+wx^{10}+x^{11}+x^{12}$, $w^{2}+w^{2}x^{2}+x^{4}+wx^{6}+w^{2}x^{8}+x^{10}+x^{12}$, $1+x+wx^{2}+w^{2}x^{3}+w^{2}x^{5}+x^{6}+wx^{7}+wx^{9}+w^{2}x^{10}+x^{11}+x^{12}$ $4$ $2\leq d_{sr}(Mat_{\mathcal{B}}(C))\leq 4$
$13$ $1+x^{13}$ $2$ $d_{sr}(Mat_{\mathcal{B}}(C))=2$
$14$ $1+x^{2}+x^{4}+x^{6}+x^{12}+x^{14}$, $1+x+x^{2}+x^{3}+x^{6}+x^{8}+x^{9}+x^{10}+x^{13}+x^{14}$, $1+x+x^{4}+x^{5}+x^{6}+x^{8}+x^{11}+x^{12}+x^{13}+x^{14}$, $1+x^{2}+x^{8}+x^{10}+x^{12}+x^{14}$ $4$ $2\leq d_{sr}(Mat_{\mathcal{B}}(C))\leq 4$
$15$ $w^{2}+wx^{2}+x^{4}+w^{2}x^{5}+wx^{7}+w^{2}x^{8}+x^{9}+x^{10}+w^{2}x^{13}+x^{15}$, $w+x+x^{2}+x^{3}+w^{2}x^{5}+x^{6}+w^{2}x^{7}+w^{2}x^{9}+wx^{12}+x^{13}+w^{2}x^{14}+x^{15}$, $1+x^{2}+w^{2}x^{4}+x^{5}+wx^{6}+x^{7}+x^{8}+w^{2}x^{9}+x^{10}+wx^{11}+x^{13}+x^{15}$, $w+w^{2}x+wx^{2}+x^{3}+w^{2}x^{6}+w^{2}x^{8}+wx^{9}+w^{2}x^{10}+wx^{12}+wx^{13}+wx^{14}+x^{15}$, $w^{2}+x^{2}+w^{2}x^{5}+w^{2}x^{6}+x^{7}+wx^{8}+x^{10}+w^{2}x^{11}+wx^{13}+x^{15}$, $1+x+x^{3}+x^{4}+x^{5}+x^{6}+x^{10}+x^{13}+x^{14}+x^{15}$, $1+x+x^{2}+x^{5}+x^{9}+x^{10}+x^{11}+x^{12}+x^{14}+x^{15}$, $w+w^{2}x^{2}+x^{4}+wx^{5}+w^{2}x^{7}+wx^{8}+x^{9}+x^{10}+wx^{13}+x^{15}$, $w^{2}+x+x^{2}+x^{3}+wx^{5}+x^{6}+wx^{7}+wx^{9}+w^{2}x^{12}+x^{13}+wx^{14}+x^{15}$, $1+x^{2}+wx^{4}+x^{5}+w^{2}x^{6}+x^{7}+x^{8}+wx^{9}+x^{10}+w^{2}x^{11}+x^{13}+x^{15}$, $w^{2}+wx+w^{2}x^{2}+x^{3}+wx^{6}+wx^{8}+w^{2}x^{9}+wx^{10}+w^{2}x^{12}+w^{2}x^{13}+w^{2}x^{14}+x^{15}$, $w+x^{2}+wx^{5}+wx^{6}+x^{7}+w^{2}x^{8}+x^{10}+wx^{11}+w^{2}x^{13}+x^{15}$ $6$ $3\leq d_{sr}(Mat_{\mathcal{B}}(C))\leq 6$

### v2

@@ 5 Asymptotically Good Self-Dual Sum-Rank Codes
In this section, we aim to prove that there exists a sequence of asymptotically good self-dual sum-rank codes. In order to construct the sequence, we will use the second method above and a sequence of asymptotically good Euclidean self-dual codes. In the proof of the following theorem, we also use an important theorem called Bolzano-Weierstrass Theorem which can be found in [49]. It claims that for any bounded sequence of real numbers, there exists a convergent subsequence of it.
Firstly, we review the sequence of asymptotically good Euclidean self-dual codes in [31]. Let $G$ be an abelian group of order $n$. By $\mathbb{F}_{q}G$ we denote the group algebra, i.e., the $\mathbb{F}_{q}$-vector space with basis $G=\{x_{1},\ldots,x_{n}\}$ and with the multiplication induced by the group multiplication of $G$. Any vector $\bm{a}=(a_{1},\ldots,a_{n})\in\mathbb{F}_{q}^{n}$ can be identified with an element $\sum\limits_{i=1}^{n}a_{i}x_{i}\in\mathbb{F}_{q}G$. An ideal $C$ of $\mathbb{F}_{q}G$ is called an $\mathbb{F}_{q}G$-code. Let $(\mathbb{F}_{q}G)^{2}:=\mathbb{F}_{q}G\times\mathbb{F}_{q}G=\{(a,b)\,|\,a,b\in\mathbb{F}_{q}G\}$ which is an $\mathbb{F}_{q}G$-module. Any $\mathbb{F}_{q}G$-submodule of $(\mathbb{F}_{q}G)^{2}$ is called a 2-quasi-$\mathbb{F}_{q}G$ code. In addition, we recall the definition of $q$-cyclotomic cosets $C_{i}^{(q,n)}=\{i,iq,\ldots,iq^{l-1}\}{\pmod{n}}$ where $l$ is the smallest positive integer such that $iq^{l}\equiv i{\pmod{n}}$. Obviously, $\{0\}$ is a $q$-cyclotomic coset which we call the trivial $q$-cyclotomic coset. By $\mu_{q}(n)$ we denote the minimal size of the non-trivial $q$-cyclotomic cosets.
Lemma 5.1.
([31, Theorem 1.1]) Let $n_{1}$, $n_{2}$, $\ldots$ be odd positive integers coprime to $q$ satisfying that $n_{i}$ goes to infinity with $i$ and $\lim\limits_{i\rightarrow\infty}\frac{\log_{q}n_{i}}{\mu_{q}(n_{i})}=0$. For $i=1,2,\ldots$, let $G_{i}$ be any abelian group of order $n_{i}$. If 1-1 is a square in $\mathbb{F}_{q}$, then there exist self-dual $2$-quasi-$\mathbb{F}_{q}G_{i}$ codes $C_{i}$, $i=1,2,\ldots$, such that the code sequence $C_{1}$, $C_{2}$, $\ldots$ is asymptotically good.
By [31], there exist odd integers $n_{1}$, $n_{2}$, $\ldots$ coprime to $q$ satisfying that $\lim\limits_{i\rightarrow\infty}n_{i}=\infty$ and $\lim\limits_{i\rightarrow\infty}\frac{\log_{q}n_{i}}{\mu_{q}(n_{i})}=0$. Hence, if 1-1 is a square in $\mathbb{F}_{q}$, there exists a sequence of asymptotically good Euclidean self-dual codes over $\mathbb{F}_{q}$. Using this sequence of asymptotically good Euclidean self-dual codes, we prove that there exist asymptotically good self-dual sum-rank codes.
@@
As $q$ is even, then 1=1-1=1 in $\mathbb{F}_{q^{2}}$. Thus, 1-1 is a square in $\mathbb{F}_{q^{2}}$. Since $t_{1}$, $t_{2}$, $\ldots$ are odd positive integers, $t_{1}$, $t_{2}$, $\ldots$ are coprime to $q^{2}=2^{2l}$. In addition, as $t_{i}$ goes to infinity with $i$ and $\lim\limits_{i\rightarrow\infty}\frac{\log_{q^{2}}t_{i}}{\mu_{q^{2}}(t_{i})}=0$, by Lemma 5.1, there exist Euclidean self-dual codes $C^{\prime}_{i}\subseteq\mathbb{F}_{q^{2}}^{2t_{i}}$, $i=1,2,\ldots$ such that the code sequence $C_{1}^{\prime}$, $C_{2}^{\prime}$, $\ldots$ is asymptotically good.
@@
As a self-dual basis $\mathcal{B}$ of $\mathbb{F}_{q}^{m}$ over $\mathbb{F}_{q}$ exists if and only if $q$ is even or both $q$ and $m$ are odd, there is no self-dual basis of $\mathbb{F}_{q^{2}}$ over $\mathbb{F}_{q}$ if $q$ is a power of an odd prime. Thus, in this case, even if $C$ is self-dual, we cannot guarantee that $Mat_{\mathcal{B}}(C)$ is a self-dual sum-rank code. This is the reason why we restrict $q$ to be even.
@@ 6 Conclusion
The target of this paper is self-dual codes and LCD codes in sum-rank metric. We introduce the notions of self-dual codes and LCD codes in sum-rank metric and give some properties of self-dual sum-rank codes. We also characterize the dual codes of two classes of sum-rank codes and provide two methods of constructing self-dual codes and LCD codes in sum-rank metric from Euclidean self-dual codes and Euclidean LCD codes. With these methods, we construct not only some normal self-dual codes and LCD codes in sum-rank metric with good parameters but also some cyclic self-dual codes and cyclic LCD codes in sum-rank metric. We also prove that asymptotically good self-dual sum-rank codes exist by the second method.
@@
    Let $C_{i}$ be a linear code in $\mathbb{F}_{q^{m}}^{t}$, $0\leq i\leq m-1$. Is it true that $SR(C_{0},C_{1},\ldots,C_{m-1})^{\bot_{tr}}=SR(C_{0}^{\bot},C_{1}^{\bot},\ldots,C_{m-1}^{\bot})$?
In addition, giving more constructions of self-dual codes and LCD codes in sum-rank metric with good parameters will be a significant direction.
@@ 88 $1+x^{4}$ 22 $d_{sr}(SR(C_{0},C_{1}))=2$
11 $1+x$ 22 $d_{sr}(Mat_{\mathcal{B}}(C))=1$
22 $1+x^{2}$ 22 $d_{sr}(Mat_{\mathcal{B}}(C))=2$
33 $^{2}+w^{2}x+x^{2}+x^{3}, w+w$$wx+x^{2}+x^{3}$ 33 $d_{sr}(Mat_{\mathcal{B}}(C))=2$
44 $1+x^{4}$ 22 $d_{sr}(Mat_{\mathcal{B}}(C))=2$
55 $1+x^{5}$ 22 $d_{sr}(Mat_{\mathcal{B}}(C))=2$
66 $^{2}+w^{2}x+x^{2}+wx^{3}+w^{2}x^{4}+x^{5}+x^{6}, w+w$$wx+x^{2}+w^{2}x^{3}+wx^{4}+x^{5}+x^{6}$ 44 $2\leq d_{sr}(Mat_{\mathcal{B}}(C))\leq 4$
77 $1+x+x^{2}+x^{3}+x^{6}+x^{7}$, $1+x+x^{4}+x^{5}+x^{6}+x^{7}$ 44 $2\leq d_{sr}(Mat_{\mathcal{B}}(C))\leq 3$
88 $1+x^{8}$ 22 $d_{sr}(Mat_{\mathcal{B}}(C))=2$
99 $^{2}+w^{2}x+x^{2}+wx^{3}+w^{2}x^{4}+x^{5}+wx^{6}+w^{2}x^{7}+x^{8}+x^{9}, w+w$$wx+x^{2}+w^{2}x^{3}+wx^{4}+x^{5}+w^{2}x^{6}+wx^{7}+x^{8}+x^{9}$ 44 $2\leq d_{sr}(Mat_{\mathcal{B}}(C))\leq 4$
$10$ $1+x^{10}$ 22 $d_{sr}(Mat_{\mathcal{B}}(C))=2$
$11$ $1+x+wx^{2}+wx^{3}+x^{4}+x^{5}+x^{6}+x^{7}+w^{2}x^{8}+w^{2}x^{9}+x^{10}+x^{11}$, $1+x+w^{2}x^{2}+w^{2}x^{3}+x^{4}+x^{5}+x^{6}+x^{7}+wx^{8}+wx^{9}+x^{10}+x^{11}$ 66 $3\leq d_{sr}(Mat_{\mathcal{B}}(C))\leq 6$
$12$ $1+x+w^{2}x^{2}+wx^{3}+wx^{5}+x^{6}+w^{2}x^{7}+w^{2}x^{9}+wx^{10}+x^{11}+x^{12}$, $w+wx^{2}+x^{4}+w^{2}x^{6}+wx^{8}+x^{10}+x^{12}$, $^{2}+w^{2}x+x^{2}+wx^{3}+w^{2}x^{4}+x^{5}+wx^{6}+w^{2}x^{7}+x^{8}+wx^{9}+w^{2}x^{10}+x^{11}+x^{12}, w+w$$wx+x^{2}+w^{2}x^{3}+wx^{4}+x^{5}+w^{2}x^{6}+wx^{7}+x^{8}+w^{2}x^{9}+wx^{10}+x^{11}+x^{12}$, $w^{2}+w^{2}x^{2}+x^{4}+wx^{6}+w^{2}x^{8}+x^{10}+x^{12}$, $1+x+wx^{2}+w^{2}x^{3}+w^{2}x^{5}+x^{6}+wx^{7}+wx^{9}+w^{2}x^{10}+x^{11}+x^{12}$ 44 $2\leq d_{sr}(Mat_{\mathcal{B}}(C))\leq 4$
$13$ $1+x^{13}$ 22 $d_{sr}(Mat_{\mathcal{B}}(C))=2$
$14$ $1+x^{2}+x^{4}+x^{6}+x^{12}+x^{14}$, $1+x+x^{2}+x^{3}+x^{6}+x^{8}+x^{9}+x^{10}+x^{13}+x^{14}$, $1+x+x^{4}+x^{5}+x^{6}+x^{8}+x^{11}+x^{12}+x^{13}+x^{14}$, $1+x^{2}+x^{8}+x^{10}+x^{12}+x^{14}$ 44 $2\leq d_{sr}(Mat_{\mathcal{B}}(C))\leq 4$
$15$ $w^{2}+wx^{2}+x^{4}+w^{2}x^{5}+wx^{7}+w^{2}x^{8}+x^{9}+x^{10}+w^{2}x^{13}+x^{15}$, $w+x+x^{2}+x^{3}+w^{2}x^{5}+x^{6}+w^{2}x^{7}+w^{2}x^{9}+wx^{12}+x^{13}+w^{2}x^{14}+x^{15}$, $1+x^{2}+w^{2}x^{4}+x^{5}+wx^{6}+x^{7}+x^{8}+w^{2}x^{9}+x^{10}+wx^{11}+x^{13}+x^{15}$, $w+w^{2}x+wx^{2}+x^{3}+w^{2}x^{6}+w^{2}x^{8}+wx^{9}+w^{2}x^{10}+wx^{12}+wx^{13}+wx^{14}+x^{15}$, $w^{2}+x^{2}+w^{2}x^{5}+w^{2}x^{6}+x^{7}+wx^{8}+x^{10}+w^{2}x^{11}+wx^{13}+x^{15}$, $1+x+x^{3}+x^{4}+x^{5}+x^{6}+x^{10}+x^{13}+x^{14}+x^{15}$, $1+x+x^{2}+x^{5}+x^{9}+x^{10}+x^{11}+x^{12}+x^{14}+x^{15}$, $w+w^{2}x^{2}+x^{4}+wx^{5}+w^{2}x^{7}+wx^{8}+x^{9}+x^{10}+wx^{13}+x^{15}$, $w^{2}+x+x^{2}+x^{3}+wx^{5}+x^{6}+wx^{7}+wx^{9}+w^{2}x^{12}+x^{13}+wx^{14}+x^{15}$, $1+x^{2}+wx^{4}+x^{5}+w^{2}x^{6}+x^{7}+x^{8}+wx^{9}+x^{10}+w^{2}x^{11}+x^{13}+x^{15}$, $w^{2}+wx+w^{2}x^{2}+x^{3}+wx^{6}+wx^{8}+w^{2}x^{9}+wx^{10}+w^{2}x^{12}+w^{2}x^{13}+w^{2}x^{14}+x^{15}$, $w+x^{2}+wx^{5}+wx^{6}+x^{7}+w^{2}x^{8}+x^{10}+wx^{11}+w^{2}x^{13}+x^{15}$ 66 $3\leq d_{sr}(Mat_{\mathcal{B}}(C))\leq 6$